OpenStudy (anonymous):

if you wanted to eliminate the x variable by adding the two equations in the following system, what could be your first step? 2x-y=6 4x-4y=10

OpenStudy (anonymous):

@rizwan_uet @qweqwe123123123123111

OpenStudy (anonymous):

You can eliminate the x variable by substituting its Y equivalent Take 2x-y=6 Isolate 2x by adding Y to both sides 2x = 6+y Get the x equivalency by dividing both sides by 2 x = 3+y/2 Now you can substitute any "x" with "3+y/2" 2x-y=6 becomes 2(3+y/2)-y=6 and 4x-4y=10 becomes 4(3+y/2)-4y=10 The "x" variable has now been eliminated.